Rampart Corporation has a dividend yield of 1.6 %. Its equity cost of capital is 7.5 %​, and its dividends are expected to grow at a constant rate.

a. What is the expected growth rate of​ Rampart's dividends?

b. What is the expected growth rate of​ Rampart's share​ price?
Answer #1

Dividend yield = D1 /PO = 1.6% Cost of equity. Ke = 7.5% Ke = D1 /PO) +g 7.5% = (1.6%) +g Growth rate, g = 7.5% - 1.6% = 5.9%